# Spatiotemporal-Augmented Graph Neural Networks for Human Mobility Simulation

### v1

@@ Abstract
Human mobility patterns have shown significant applications in policy-decision scenarios and economic behavior researches. The human mobility simulation task aims to generate human mobility trajectories given a small set of trajectory data, which have aroused much concern due to the scarcity and sparsity of human mobility data. Existing methods mostly rely on the static relationships of locations, while largely neglect the dynamic spatiotemporal effects of locations. On the one hand, spatiotemporal correspondences of visit distributions reveal the spatial proximity and the functionality similarity of locations. On the other hand, the varying durations in different locations hinder the iterative generation process of the mobility trajectory. Therefore, we propose a novel framework to model the dynamic spatiotemporal effects of locations, namely SpatioTemporal-Augmented gRaph neural networks (STAR). The STAR framework designs various spatiotemporal graphs to capture the spatiotemporal correspondences and builds a novel dwell branch to simulate the varying durations in locations, which is finally optimized in an adversarial manner. The comprehensive experiments over four real datasets for the human mobility simulation have verified the superiority of STAR to state-of-the-art methods. Our code will be made publicly available.
@@ 1 Introduction
Human mobility patterns have attracted widespread attention [1, 2, 3] to investigate when and where people’s activities happen, and aroused multi-disciplinary applications such as urban planning [4], pollution abatement [5], and epidemic prevention [6]. For example, reliable human mobility models can inform policy-making of crucial projects, such as planning for expanding urban land [7] and making effective controlling strategies of gross polluters [5]. Specifically, the human mobility patterns have been investigated to deal with the COVID-19 pandemic, appropriately informing the re-opening policys [8] and allocating the limited medical resources [6].
@@
Figure 1: An illustration example. (a) depicts that various locations (HOME, Bar, Cafe, Pizza) get busy when people get off work from the office, revealing the spatiotemporal correspondences among locations. (b) depicts the successive visits of locations in an individual trajectory, resulting varying durations in different locations.
@@
However, it is difficult to obtain human mobility trajectories recording the spatiotemporal sequences of human mobility locations due to privacy concerns and commercial limitations. The human mobility simulation task aims to generate massive artificial human mobility trajectories with high fidelity given a small set of real-world human mobility data, which can largely alleviate the scarcity and sparsity of existing data. Moreover, such controllable generations of human mobility data can perform counterfactual inference to measure the treatment effects of different strategies [8] and investigate the emergence patterns of urban growth when increasing the scales of human mobility behaviors [9, 4].
@@
Capturing the regularities of human mobility behaviors lies at the core of human mobility simulation. Previous model-based methods [10, 11] build various conditional transition matrices of locations and maximize the log-likelihood by fitting the real-world data. Despite the simplicity and interpretability, these methods rely on large-scale mobility trajectory data and fine-grained labels of location categories, which limits their practical utility. Recently, deep learning-based methods [12, 13, 14] introduce the static relationships of locations to generate human mobility sequences in a global view. DeepMove [12] captures the sequential dependency within a single trajectory with the neural attention layer; CGE [14] obtains the global relationships of locations by the proposed static relationship matrix of locations; MoveSim [13] takes a step further to introduce the structure prior of locations with an attention layer. However, these methods focus on the static relationships of locations, while the dynamic spatiotemporal effects of locations are largely under-explored.
@@
As shown in Figure 1, the spatiotemporal effects of locations can be observed from two aspects: the spatiotemporal correspondences and the varying durations. On the one hand, Figure 1 (a) depicts that various locations (Home, Bar, Cafe, Pizza) get busy when people get off work. It indicates the spatiotemporal correspondences between these locations, which can reflect both the spatial proximity and the functionality similarity. On the other hand, Figure 1 (b) depicts the varying durations in different locations of an illustrative trajectory, which describes the spatiotemporal continuity of human mobility behaviors. Existing methods generate the simulation trajectories without considering the varying durations, where locations with short dwell time will undoubtedly get neglected in the optimization goals.
@@
  • A novel framework STAR is proposed to handle the spatiotemporal correspondences and the varying durations with the multi-channel embedding module and the dual-branch decision generator module, respectively.
  • Extensive experiments on various real-world datasets demonstrate that the proposed STAR consistently outperforms the state-of-the-art baselines in human mobility simulation with high fidelity. The ablation studies further reveal the working mechanisms of the spatiotemporal graphs and the dwell branch.
@@ 2.1 Human Mobility Simulation
The human mobility simulation task aims to generate artificial mobility trajectories with realistic mobility patterns given a small set of human mobility data [16, 17, 18, 19, 13, 20]. The generated artificial trajectories must reproduce a set of spatial and temporal mobility patterns, such as the distribution of characteristic distances and the predictability of human whereabouts. Temporal patterns usually include the number and sequence of visited locations together with the time and duration of the visits, which involves balancing an individual’s routine and sporadic out-of-routine mobility patterns. And spatial patterns include the preference for short distances[1, 21], the tendency to split into returnees and explorers[22], and the fact of visiting multiple sites for constant times for individuals[23].
@@
In the early stage, Markov-based models dominate the task of human mobility simulation. For example, First-order MC [24] defines the state as the location accessed and assumes that the next location depends only on the current location, thus constructing a transition matrix to capture the first-order transition probability among locations. HMM [25] is established with discrete emission probability and optimized by the Baum-Welch algorithm. IO-HMM [11] further extends Markov-based models by introducing more annotation information, which also improves interpretability. However, Markov-based models are limited in capturing long-term dependencies and incorporating individual preferences.
@@
To make up the deficiency of Markov-based models, a large body of mechanistic methods have emerged, which can reproduce basic temporal, spatial, and social patterns of human mobility [19, 26, 16, 17, 18]. For example, in the Exploration and Preferential Return (EPR) model[27], an agent can select a new location that has never been visited before based on a random walk process utilizing a power-law jump-size distribution, or return to a previously visited location based on its frequency. Then several studies enhanced the EPR model by incorporating increasingly elaborate spatial or social mechanisms [22, 28, 23, 29, 30]. EPR and its extensions primarily focus on the spatial patterns of human mobility, neglecting to capture the temporal mechanisms. TimeGeo [31] and DITRAS [26] improve temporal mechanism by integrating a data-driven model into an EPR-like model to capture both routine and out-of-routine circadian preferences. Despite the interpretability of mechanistic methods, their realism is limited by the simplicity of the implemented mechanisms.
@@
The limitations mentioned above can be addressed by deep learning generative paradigms such as recurrent neural networks (RNNs) and generative adversarial networks (GANs), which can learn the distribution of data by capturing complex and non-linear relationships in data and generate mobility trajectories from the same distribution. As a result, the deep learning approaches for human mobility simulation can generate more realistic data than traditional methods. RNN-based models prefer to maximize the prediction likelihood of the next location likelihood [32, 33], resulting in ignoring the long-term influence and the so-called exposure bias. SeqGAN [33], the pioneering work of sequence generation based on Generative Adversarial Networks (GAN) [15], proposes to optimize the generator of discrete tokens by a policy gradient technique. Based on SeqGAN, MoveSim [13] further introduces the location structure as a prior for human mobility simulation. ActSTD [34] improves the dynamic modeling of individual trajectories by the neural ordinary equation. It is worth noting that the purpose of human mobility simulation differs from that of human mobility prediction [35, 36, 37]. The former aims to produce results that reflect the characteristics of real-world data, which should not be too similar to real data so as to protect user privacy. Conversely, the latter emphasizes testing the model’s ability to recover the real data. Despite deep learning approaches proposed for human mobility prediction [38, 39], simulating daily mobility has been underexplored.
@@ 2.2 Graph Neural Networks
Recently, the rapid development of deep learning [40, 41] has inspired the researches on Graph Neural Networks (GNNs) [42, 43, 44, 45], which attracts much concern from various fields due to their widespread applications for non-grid data. Graph Convolution Networks (GCN) [43] advances the node classification task by a large margin compared to the previous methods, showing a pioneering attempt at GNNs; GraphSAGE [44] aims to perform graph learning on large-scale datasets by sampling the recursively expanded ego-graphs; Graph Attention Network (GAT) [45] weights the neighborhood importances by an adaptive attention mechanism [46]. Due to their significant progress, GNNs have been applied in various fields such as social networks, molecular structures, traffic flows, and so on.
@@
Among the numerous fields in which GNNs are applied, urban computing is most relevant to the human mobility simulation task in this paper. Urban computing aims to understand the urban patterns and dynamics from different application domains where the big data explodes, such as transportation, environment, security, etc [47]. Due to the spatio-temporal characteristics of some typical urban data, such as traffic network flow [48, 49, 50, 51], crowd flow [52], environmental monitoring data, etc., some previous works combines graph neural networks with various temporal learning networks to capture the dynamics in the spacial and temporal dimensions [53]. The hybrid neural network architecture is collectively referred to as spatio-temporal graph neural network (STGNN).
@@
The basic STGNN framework for predictive learning is composed of three modules: Data Processing Module (DPM) which constructs the spatio-temporal graph from raw data, Spatio-Temporal Graph Learning Module (STGLM) which extracts hidden spatio-temporal dependencies within complex social systems and Task-Aware Prediction Module (TPM) which maps the spatio-temporal hidden representation from STGLM into the space of downstream prediction tasks. As the most crucial part in STGNN, STGLM combines spatial learning networks such as spectral graph convolutional networks (Spectral GCNs), spatial graph convolutional networks (Spatial GCNs) or GATs, and temporal leaning networks such as RNNs, temporal convolutional networks (TCNs) or temporal self-attention networks (TSANs) organically through a certain spatio-temporal fusion neural architecture.
@@
Therefore, almost all current researches focus on the design of the neural architectures in STGLM and there are many frontier methods to improve the learning of spatio-temporal dependencies. For example, THINK [54] and DMGCRN [55] perform the hyperbolic graph neural network on the Poincare ball to directly capture multi-scale spatial dependencies. ASTGCN [56] employ a typical three-branch architecture for multi-granularity temporal learning, where the data undergone the calculation of multiple GCNs and Attention networks from the three branches will be finally fused by the learnable weight matrix. STSGCN [57] fuses spatio-temporal dependencies by constructing the spatio-temporal synchronous graph. Based on STSGNN, STFGNN [58] introduces topology-based graph and similarity-based graph simultaneously to construct a spatio-temporal synchronous graph, making the spatio-temporal synchronous graph more informative. S2TAT [59] proposes a spatio-temporal synchronous transformer framework to enhance the learning capability with attention mechanisms.
@@
However, there are two main differences between our work and the existing researches on STGNN. First, in previous studies, the topological structure of the spatio-temporal graph (e.g. road network) is fixed, while ours is self-constructed and closely related to the order of the locations visited by individuals. Second, the existing methods based on STGNN are almost used for predictive learning tasks in urban computing, but our approach focuses on simulation task which emphasizes the effective capture of the overall patterns rather than the accurate prediction of a single entity.
@@ 3 Problem Statement
A human mobility trajectory can be defined as a spatiotemporal sequence $s=[\tau_{1},\tau_{2},\cdots,\tau_{L}]$. The $l$-th visit record $\tau_{l}$ represents a tuple $(p_{l},t_{l})$, where $p_{l}$ records the location ID and $t_{l}$ records the visit timestamp. The human mobility simulation problem is thus defined as follows.
DEFINITION 1. (Human Mobility Simulation). Given a real-world mobility trajectory dataset $\mathcal{S}=\{s_{1},s_{2},\cdots,s_{m}\}$, our goal is learning to simulate human mobility behaviors in order to generate an artificial trajectory $\hat{s}=\left[\hat{\tau}_{1},\hat{\tau}_{2},\dots,\hat{\tau}_{L}\right]$ with fidelity and utility.
Directly maximizing the likelihood of the sequence generation model would undoubtedly cause the eposure bias [32, 33] towards the training data, resulting in poor generalization abilities. Therefore, it is useful to advance the human mobility simulation under the framework of SeqGAN [15, 33], which coordinates the optimization of the generator (sequence generation model) and a discriminator in a min-max game, written as
@@ 4 Methods
The proposed SpatioTemporal-Augmented gRaph neural networks (STAR) aims to capture the spatiotemporal correspondences among locations and the varying durations in different locations to advance the human mobility simulation task in a spatiotemporal way. As illustrated in Figure 2, the STAR framework consists of three key components: (i) the multi-channel embedding module learns spatiotemporal semantics of locations by the dense representations based on the proposed multi-channel location graphs with the partial observed human trajectories; (ii) the decision generator module tackles the varying duration problem by the exploration branch and the dwell branch, alleviating the effects of highly repetitive patterns in mobility trajectories; (iii) the policy discriminator module provides the rewards for the generator at each step based on the policy gradient technique.

### v3

@@ Abstract
Human mobility patterns have shown significant applications in policy-decision scenarios and economic behavior researches. The human mobility simulation task aims to generate human mobility trajectories given a small set of trajectory data, which have aroused much concern due to the scarcity and sparsity of human mobility data. Existing methods mostly rely on the static relationships of locations, while largely neglect the dynamic spatiotemporal effects of locations. On the one hand, spatiotemporal correspondences of visit distributions reveal the spatial proximity and the functionality similarity of locations. On the other hand, the varying durations in different locations hinder the iterative generation process of the mobility trajectory. Therefore, we propose a novel framework to model the dynamic spatiotemporal effects of locations, namely SpatioTemporal-Augmented gRaph neural networks (STAR). The STAR framework designs various spatiotemporal graphs to capture the spatiotemporal correspondences and builds a novel dwell branch to simulate the varying durations in locations, which is finally optimized in an adversarial manner. The comprehensive experiments over four real datasets for the human mobility simulation have verified the superiority of STAR to state-of-the-art methods. Our code is available at https://github.com/Star607/STAR-TKDE.
@@ 1 Introduction
Human mobility patterns have attracted widespread attention [1, 2, 3] to investigate when and where people’s activities happen, and aroused multi-disciplinary applications such as urban planning [4], pollution abatement [5], and epidemic prevention [6]. For example, reliable human mobility models can inform policy-making of crucial projects, such as planning for expanding urban land [7] and making effective controlling strategies of gross polluters [5]. Specifically, the human mobility patterns have been investigated to deal with the COVID-19 pandemic, appropriately informing the reopening policies [8] and allocating the limited medical resources [6].
@@
Figure 1: An illustration example. (a) depicts that various locations (HOME, Bar, Cafe, Pizza) get busy when people get off work from the office, revealing the spatiotemporal correspondences among locations. (b) depicts the successive visits of locations in an individual trajectory, resulting in varying durations in different locations.
@@
However, it is difficult to obtain human mobility trajectories recording the spatiotemporal sequences of human mobility locations due to privacy concerns and commercial limitations. The human mobility simulation task aims to generate massive artificial human mobility trajectories with high fidelity given a small set of real-world human mobility data, which can largely alleviate the scarcity and sparsity of existing data. Such controllable generations of human mobility data can perform counterfactual inference to measure the treatment effects of different strategies [8] and investigate the emergence patterns of urban growth when increasing human mobility scales [9, 4].
@@
Capturing the regularities of human mobility behaviors lies at the core of human mobility simulation. Previous model-based methods [10, 11] build various conditional transition matrices of locations and maximize the log-likelihood by fitting the real-world data. Despite the simplicity and interpretability, these methods rely on large-scale mobility trajectory data and fine-grained labels of location categories, which limits their practical utility. Recently, deep learning-based methods [12, 13, 14] introduce the static relationships of locations to generate human mobility sequences in a global view. DeepMove [12] captures the sequential dependency within a single trajectory with the neural attention layer; CGE [14] obtains the global relationships of locations by the proposed static relationship matrix of locations; MoveSim [13] takes a step further to introduce the structure prior of locations with an attention layer. However, these methods focus on the static relationships of locations, while the dynamic spatiotemporal effects of locations are largely underexplored.
@@
As shown in Figure 1, the spatiotemporal effects of locations can be observed from two aspects: the spatiotemporal correspondences and the varying durations. On the one hand, Figure 1 (a) depicts that various locations (Home, Bar, Cafe, Pizza) get busy when people get off work. It indicates spatiotemporal correspondences between these locations, which can reflect both the spatial proximity and the functionality similarity. On the other hand, Figure 1 (b) depicts the varying durations in different locations of an illustrative trajectory, which describes the spatiotemporal continuity of human mobility behaviors. Existing methods generate the simulation trajectories without considering the varying durations, where locations with short dwell time will undoubtedly get neglected in the optimization goals.
@@
  • A novel framework STAR is proposed to handle spatiotemporal correspondences and varying durations with the multi-channel embedding module and the dual-branch decision generator module, respectively.
  • Extensive experiments on various datasets demonstrate that our proposed STAR consistently outperforms the state-of-the-art baselines in human mobility simulation with high fidelity. The ablation studies further reveal the working mechanisms of the spatiotemporal graphs and the dwell branch.
@@ 2.1 Human Mobility Simulation
The human mobility simulation task aims to generate artificial mobility trajectories with realistic mobility patterns given a small set of human mobility data [16, 17, 18, 19, 20]. The generated artificial trajectories must reproduce a set of spatial and temporal mobility patterns, such as the distribution of characteristic distances and the predictability of human whereabouts. Temporal patterns usually include the number and sequence of visited locations together with the time and duration of the visits, which involves balancing an individual’s routine and sporadic out-of-routine mobility patterns. And spatial patterns include the preference for short distances[1, 21], the tendency to split into returnees and explorers[22], and the fact of visiting multiple sites for constant times for individuals[23].
@@
In the early stages, Markov-based models dominate the human mobility simulation task. For example, the first-order MC [24] defines the state as the accessed location and assumes that the next location solely depends on the current location, thus constructing a transition matrix to capture the first-order transition probability among locations. HMM [25] is established with discrete emission probability and optimized by the Baum-Welch algorithm. IO-HMM [11] further extends HMM by introducing more annotation information, which also improves interpretability. However, Markov-based models are limited in capturing long-term dependencies and incorporating individual preferences.
@@
To make up for the deficiency of Markov-based models, a large body of mechanistic methods have emerged, which can reproduce basic temporal, spatial, and social patterns of human mobility [26, 27, 16, 17, 28]. For example, Exploration and Preferential Return (EPR) model[29] enables an agent to select a new location that has never been visited before based on a random walk process utilizing a power-law jump-size distribution, or return to a previously visited location based on its frequency. Then several studies enhanced the EPR model by incorporating increasingly elaborate spatial or social mechanisms [22, 30, 23, 31, 32]. EPR and its extensions primarily focus on capturing the spatial patterns of human mobility, neglecting to capture the temporal mechanisms. TimeGeo [33] and DITRAS [27] improve the temporal mechanism by integrating a data-driven model into an EPR-like model to capture both routine and out-of-routine circadian preferences. Despite the interpretability of mechanistic methods, their realism is limited by the simplicity of the implemented mechanisms.
@@
The limitations mentioned above can be addressed by deep learning generative paradigms such as recurrent neural networks (RNNs) and generative adversarial networks (GANs), which can learn the distribution of data by capturing complex and non-linear relationships in data and generate mobility trajectories from the same distribution. RNN-based models prefer to maximize the prediction likelihood of the next location likelihood [34, 35], resulting in ignoring the long-term influence and the so-called exposure bias. SeqGAN [35] is the pioneering work of sequence generation based on GAN [15]. MoveSim [13] extends SeqGAN by introducing the location structure as a prior for human mobility simulation. ActSTD [36] improves the dynamic modeling of individual trajectories by the neural ordinary equation. SAND [37] combines generative adversarial imitation learning with Maslow’s need theory to simulate human activities. It is worth noting that human mobility simulation is distinct from time series prediction based on mobility [38, 39], where the former aims to generate a sequence of discrete locations while the latter focuses on predicting a real number. Furthermore, different from human mobility prediction [40, 41] and location recommendation [42], human mobility simulation emphasizes producing trajectories that reflect the characteristics of real-world data while protecting user privacy, rather than testing the model’s ability to recover the real data. Despite deep learning approaches proposed for human mobility prediction, simulating daily mobility has been underexplored.
@@ 2.2 Graph Neural Networks
Recently, the rapid development of deep learning [43, 44] has inspired the researches on Graph Neural Networks (GNNs) [45, 46, 47, 48]. Due to the significant progress of GNNs, they have been applied in various fields, where urban computing is most relevant to the human mobility simulation task in this paper. Urban computing aims to understand the urban patterns and dynamics from different application domains where the big data explodes, such as transportation, environment, security, etc [49]. Due to the spatio-temporal characteristics of some typical urban data, such as traffic network flow [50, 51, 52, 53], crowd flow [54], environmental monitoring data, etc., some previous works combine graph neural networks with various temporal learning networks to capture the dynamics in the spatial and temporal dimensions [55]. The hybrid neural network architecture is collectively referred to as spatio-temporal graph neural network (STGNN).
@@
The basic STGNN framework for predictive learning is composed of three modules: Data Processing Module (DPM) which constructs the spatio-temporal graph from raw data, Spatio-Temporal Graph Learning Module (STGLM) which extracts hidden spatio-temporal dependencies within complex social systems and Task-Aware Prediction Module (TPM) which maps the spatio-temporal hidden representation from STGLM into the space of downstream prediction tasks. As the most crucial part in STGNN, STGLM combines spatial learning networks such as spectral GCNs, spatial GCNs or GATs, and temporal learning networks such as RNNs, temporal convolutional networks (TCNs) or temporal self-attention networks (TSANs) organically through a certain spatio-temporal fusion neural architecture.
@@
Therefore, almost all current researches focus on the design of the neural architectures in STGLM and there are many frontier methods to improve the learning of spatio-temporal dependencies. For example, THINK [56] and DMGCRN [57] perform the hyperbolic graph neural network on the Poincare ball to directly capture multi-scale spatial dependencies. ASTGCN [58] employ a typical three-branch architecture for multi-granularity temporal learning, where the data undergoes calculations of multiple GCNs and Attention networks from the three branches will be finally fused using the learnable weight matrix. STSGCN [59] fuses spatio-temporal dependencies by constructing the spatio-temporal synchronous graph. Based on STSGNN, STFGNN [60] introduces both topology-based graph and similarity-based graph to construct a spatio-temporal synchronous graph, making the spatio-temporal synchronous graph more informative. S2TAT [61] proposes a spatio-temporal synchronous transformer framework to enhance the learning capability with attention mechanisms.
@@
However, there are two main differences between our work and the existing researches on STGNN [58, 59, 60, 57, 56, 61, 62]. First, the topological structure of the spatio-temporal graph (e.g. road network) in previous studies is fixed, while ours is self-constructed and closely related to the order of locations visited by individuals. Second, the existing methods based on STGNN are almost used for predictive learning tasks in urban computing, but our approach focuses on the simulation task which emphasizes the effective capture of the overall patterns rather than the accurate prediction of a single entity.
@@ 3 Problem Statement
A human mobility trajectory can be defined as a spatiotemporal sequence $s=[\tau_{1},\tau_{2},\cdots,\tau_{L}]$. The $l$-th visit record $\tau_{l}$ represents a tuple $(p_{l},t_{l})$, where $p_{l}$ records the location ID and $t_{l}$ records the visiting timestamp. The human mobility simulation problem is thus defined as follows.
DEFINITION 1. (Human Mobility Simulation). Given a real-world mobility trajectory dataset $\mathcal{S}=\{s_{1},s_{2},\cdots,s_{m}\}$, our goal is to learn to simulate human mobility behaviors in order to generate an artificial trajectory $\hat{s}=\left[\hat{\tau}_{1},\hat{\tau}_{2},\dots,\hat{\tau}_{L}\right]$ with fidelity and utility.
Directly maximizing the likelihood of the sequence generation model would undoubtedly cause the exposure bias [34, 35] towards the training data, resulting in poor generalization abilities. Therefore, it is useful to advance the human mobility simulation under the framework of SeqGAN [15, 35], which coordinates the optimization of the generator (sequence generation model) and a discriminator in a min-max game, written as
@@ 4 Methods
In this section, we propose SpatioTemporal-Augmented gRaph neural networks (STAR) to simulate human mobility, whose framework is illustrated in Figure 2. In this framework, we firstly develop the multi-channel embedding module to capture the spatiotemporal correspondences among locations based on the proposed multi-channel location graphs. Secondly, we design the decision generator module involving exploration and dwell branches to learn the varying durations in different locations from trajectories with highly repetitive patterns. Thirdly, we optimize the training process using the policy discriminator module, which provides the rewards for the generator at each step based on the policy gradient technique.
@@ 4.1 Multi-channel Embedding Module
It’s fundamental to acquire high-quality representations of locations with ample spatiotemporal semantics for effective human mobility simulation. Considering the non-grid structure of locations, we leverage GNNs to model the spatiotemporal dependencies among the locations. However, relying solely on static spatial distance or temporal transition, or both, is insufficient to capture the intricate spatiotemporal relationships. Therefore, to go beyond the geographical effects, we first build a Spatial Distance Graph (SDG) and Temporal Transition Graph (TTG) based on the observed human mobility trajectories, which are constructed from the spatial and temporal perspectives, respectively. Secondly, we propose a SpatioTemporal Graph (STG) by measuring the Wasserstein distance [63] of location visit distributions to capture the spatiotemporal effects, which combines the spatial and temporal dynamics of locations simultaneously. Finally, we fuse the representations of locations from these three graphs to obtain comprehensive semantics. In this part, we first describe the details of constructing location graphs, followed by an explanation of how location embeddings are generated.